determine the area of the region indicated.

1) -1 < Z <1

2) -2 < Z < 2



1
Expert's answer
2021-10-20T04:22:48-0400

1)

"P(-1<z<1)=0.84134-0.15866=0.68268"


2)

"P(-2<z<2)=0.97725-0.02275=0.9545"